The information that follows pertains to Richards Refrigeration, Inc.:

  1. At December 31, 2018, temporary differences existed between the financial statement carrying amounts and the tax bases of the following:
($ in millions)
Carrying
Amount
Tax
Basis
Future Taxable
(Deductible)
Amount
Buildings and equipment (net of accumulated depreciation) $ 132 $ 96 $ 36
Prepaid insurance 56 0 56
Liability—loss contingency 31 0 (31 )
  1. No temporary differences existed at the beginning of 2018.
  2. Pretax accounting income was $206 million and taxable income was $145 million for the year ended December 31, 2018. The tax rate is 30%.

Required:
1. Complete the following table given below and prepare the appropriate journal entry to record income taxes for 2018
2. What is the 2018 net income?

Complete the following table given below to record income taxes for 2018. (Enter your answers in millions rounded to 1 decimal place (i.e., 5,500,000 should be entered as 5.5). Negative amounts should be entered with a minus sign.)

($ in millions) x Tax Rate % = Tax $ Recorded as:
Pretax accounting income $206.0
Permanent differences
Income subject to taxation 206.0 x = $0.0
Temporary Differences
Depreciation x =
Prepaid insurance x =
Liability - loss contingency x =
Income taxable in current year $206.0 x =

1. 1 GJ Record 2018 income taxes.

required 2

What is the 2018 net income? (Enter your answers in millions rounded to 1 decimal place (i.e., 5,500,000 should be entered as 5.5).)

Net income million
Answer #1

ANSWER

Part-1 $ in Million * Tax Rate % = Tax $ Recorded as
Pretax accounting Income 206
Permanent Difference 0
Income Subject to taxation 206 * 30% = 61.8 Income Tax Expense
Temporary Difference:
Depreciation -36 * 30% = 10.8 Deferred Tax Liability
Prepaid Insurance -56 * 30% = 16.8 Deferred Tax Liability
Liability-Contingent Loss 31 * 30% = 9.3 Deferred Tax Asset
Income Taxable in current year 145 * 30% = 43.5 Income Tax Payable
Journal Entry:
Debit Credit
Income Tax Expense 61.8
Deferred Tax Asset 9.3
Deferred Tax Liability

27.6

Income Tax Payable 43.5
Part-2 2018 Net Income
Pretax Income 206
Less: Income Tax 30% 61.8
Net Income 144.2

_____________________________________________

If you have any query or any Explanation please ask me in the comment box, i am here to helps you.please give me positive rating.

*****************THANK YOU**************